OpenStudy (anonymous):

find the product 2x(-2x-3) pleeeease<3

OpenStudy (anonymous):

-4x - 6x

OpenStudy (anonymous):

I mean -4x squared - 6x

OpenStudy (anonymous):

its like the next level of your last question :)

OpenStudy (anonymous):

i knooooooow, like i get it if someone breaks it down and helps me through it then i look at the next one, and i'm lost. =(

OpenStudy (anonymous):

i dont know anything, lol.

OpenStudy (anonymous):

no, but you can always learn... ill break this one down into a sentence " two ex times negative two ex minus two ex times three" or (2x times -2x) - (2x times -3)

OpenStudy (anonymous):

so 2x times -2x is -4x²

OpenStudy (anonymous):

oh goooodness, i get it! :)

OpenStudy (anonymous):

woot, you're great.

OpenStudy (anonymous):

can i ask you something (about the maths) ? just to make sure you understand it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

4b(-5b+6) solve for b

OpenStudy (anonymous):

i have no idea. D: i was just taking a chance and i got 4b lol. i really dont get math. but if you explain it like you did the one before, i'd get it.

OpenStudy (anonymous):

actually im just doing that one i made up now, and its a bit harder than i thought try 1b(2b+3) solve for b?

OpenStudy (anonymous):

5b? i dont know. :/

OpenStudy (anonymous):

sory made another hard one we will make it solve for b where 0=3(2b+3)

OpenStudy (anonymous):

i know that can be done

OpenStudy (anonymous):

lol i dont know how!

OpenStudy (anonymous):

if i say expand it can you do that?

OpenStudy (anonymous):

expand it is the same as 'multiply it out' meaning 3(2b+3) = 6b + 9 = 0

OpenStudy (anonymous):

then you can drop the far left and have 6b + 9 = 0 then go 6b = -9 then go b= -9/6 hopefully that helps seeing as you moved onto another question already

OpenStudy (anonymous):

i really dont get this stuff.. lol but you should explain -4(v+1) for me, like write it out like you did helping me with the other one.=)